# RoboTransfer: Geometry-Consistent Video Diffusion for Robotic Visual Policy Transfer

### v1

@@ A.3 Appearance conditions
Objects and Background conditions:To extract the object conditions, we first segment image patches corresponding to individual objects from the full image. Each patch is resized to 384×384 and passed through the CLIP model [47] to obtain RGB embeddings. For the background conditions, we mask out all detected objects from the full image using their corresponding masks. The masked image is then fed into an inpainting model to reconstruct the background with the objects removed.
Appendix B Visual Policy Implementation Details
@@
B.1 Data Collection and Preprocessing
To ensure the validity of our evaluation, we excluded the dataset described in Section A from the real-robot experiments. The data preparation proceeded as follows:
Real Expert Data: We collected 100 sets of expert demonstration data for each manipulation task using the ALOHA teleoperation system. The primary observations were RGB images (1280x720 resolution, subsequently downscaled to 640$\times$360 to optimize training), with camera streams captured at approximately 30Hz and sampled at 10Hz for policy training. Auxiliary robot state information (e.g., joint positions, end-effector poses) was originally recorded at 200Hz and downsampled to 50Hz for model use.
Synthetic Data Generation: To further enhance model performance and generalization, we generated synthetic data based on the collected real data, specifically targeting variations in foreground objects and background scenes, for use in the subsequent fine-tuning phase.
B.2 Training Pipeline and Core Parameters
Our training pipeline is based on the ACT (Action Chunking Transformer) architecture, which processes visual input from three distinct perspectives (two wrist-mounted cameras and one fixed overhead camera). At each training timestep, the model receives a single RGB frame from each view.
Training Objective and Two-Stage Training Strategy: The policy model is designed to predict a sequence of the next 100 robot states, corresponding to a 2-second action horizon given the 50Hz robot state frequency.
We adopted a "pre-train-then-finetune" paradigm:
Pretraining on Real Data:The model was first pretrained for 100k steps using the collected real expert demonstration data. During this phase, the batch size was set to 512, and the learning rate was $1\times 10^{-4}$.
Finetuning with Synthetic Data: After pretraining, the aforementioned synthetic data was introduced to fine-tune the model for an additional 50k steps. The learning rate for this phase was reduced to $1\times 10^{-5}$.
@@ B.2 Training Pipeline and Core Parameters
B.3 Real-Robot Deployment and Evaluation
The trained policies were deployed and evaluated in the following environment:
@@ B.3 Real-Robot Deployment and Evaluation
Deployment Platform: Experiments were conducted on the Agilex Cobot Magic platform8, the same platform used for initial data collection, ensuring consistency between training and testing environments.
@@
Inference Procedure: During deployment, the policy operated in a synchronous execution mode. Specifically, the system waits for the robot to fully execute the action sequence (comprising 100 actions) generated from the model’s previous inference step before capturing new observations and proceeding to the next model inference and subsequent execution. At each decision point, the model receives as input a single RGB frame from each of its multiple camera views and the current robot state and outputs the sequence of 100 actions. The latency for this single model inference process is approximately 10ms. These actions are then executed on the robot at a 50Hz control frequency, consistent with the robot state prediction rate during training.
Appendix C More Qualitative Results
C.1 Real-to-Real Transfer
The proposed RoboTransfer offers strong controllability over both foreground and background generation. Given the same structured input, the model allows for flexible editing of background attributes such as texture and color, as shown in Figure 12. Meanwhile, as illustrated in Figure 13, the appearance of foreground objects, including their color, can also be effectively modified. This real-to-real generation framework enriches the diversity of training data for downstream tasks, contributing to better generalization of the policy model.
[…]
C.2 Sim-to-Real Transfer
As shown in Figure 14, RoboTransfer generates photorealistic videos from simulated structural inputs, including out-of-distribution cases. This Sim-to-Real paradigm reduces dependency on structured information from real-world datasets, thereby better supporting downstream robotic learning tasks.
[…]
Figure 14: Visualizations of Sim-to-Real generation results.
C.3 Comparison
Figure 15 illustrates that Cosmos Transfer [2] is mainly effective for robot scene generation under fixed camera settings (central view), while its quality drops notably for dynamic camera viewpoints (side views). Compared to Cosmos Transfer [2], RoboTransfer demonstrates significantly better multi-view consistency, enabling more realistic and coherent novel-view synthesis across varying camera poses.
[…]
Figure 15: Comparison with Cosmos generation results.

### v2

@@ 8.3 Appearance Conditions Construction
Objects and Background Conditions. Individual object patches are resized to $224\times 224$ and passed through CLIP [40] to obtain embeddings. Background conditions are generated by masking out all detected objects and applying inpainting to reconstruct the object-free scene.
9 Robot Policy Model Implementation Details
@@
9.1 Data Collection and Preprocessing
To ensure a fair evaluation, we excluded the RoboTransfer training dataset (Sec. 8.1) from the real-robot experiments. Data preparation was conducted as follows:
Real Expert Data We collected 100 expert demonstration sets per manipulation task using the ALOHA teleoperation system. Observations included RGB images at 1280×720 resolution, downscaled to 640×360 for training efficiency, captured at 30Hz, and sampled at 10Hz. Auxiliary robot state information, including joint positions and end-effector poses, was recorded at 200Hz and downsampled to 50Hz for policy training.
Synthetic Data Generation for Policy Fine-tuning. To improve generalization, we generated synthetic videos based on the real demonstrations, introducing variations in foreground objects and background scenes. The pretrained diffusion model conditioned each synthetic video on: (1) per-frame 3D geometry inputs (depth and normal maps) from real demonstrations, and (2) a reference image containing novel objects and backgrounds from a held-out set independent of the policy training data.
9.2 Training Pipeline and Core Parameters
Our training pipeline uses the ACT (Action Chunking Transformer [60]) architecture, processing visual input from three cameras (two wrist-mounted, one overhead). At each timestep, the model receives one RGB frame per view.
Training Objective and Strategy. The policy predicts the next 100 robot states (2s horizon at 50Hz). We adopt a pretrain-then-finetune strategy: 1) Pretraining on Real Data: The model was first pretrained for 100k steps using the collected real expert demonstration data. During this phase, the batch size was set to 512, and the learning rate was $1\times 10^{-4}$. 2) Finetuning with Synthetic Data: After pretraining, the synthetic data was introduced to fine-tune the model for an additional 50k steps. The learning rate for this phase was reduced to $1\times 10^{-5}$.
@@ 9.2 Training Pipeline and Core Parameters
9.3 Real-Robot Deployment and Evaluation
@@
Deployment Platform. Policies were evaluated on the Agilex Cobot Magic platform (Figure 12), the same system used for data collection to ensure consistency between training and evaluation environments.
@@
Inference Procedure During deployment, the policy operates synchronously: the robot executes the full action sequence (100 actions) generated from the previous inference step before capturing new observations. At each decision point, the model receives one RGB frame per camera view along with the current robot state and outputs the next 100 actions. Inference latency is  10ms, and actions are executed at 50Hz, matching the training robot state frequency.
10 Discussion
10.1 Analysis on Geometry-Appearance Decoupling
In our framework, reference images serve primarily as appearance conditions, capturing visual properties like texture and color, rather than imposing strict geometric constraints. Global 3D geometry conditions ensure multi-view and temporal consistency across frames.
1. Geometry conditions. Multi-view depth and normal maps encode precise 3D structure and scene layout, even under dynamic wrist-mounted camera motion, defining where and how content should be rendered.
2. Appearance transfer. The reference image provides the target visual style (e.g., textures or background appearance) that the diffusion model transfers onto the global geometry. The reference does not need to be aligned with each frame; the model synthesizes a consistent appearance across views and time by using geometry as an anchor.
Qualitative results show visually coherent, realistic videos in wrist-view settings. Real-to-Real experiments confirm effective appearance transfer to dynamic robotic scenes, while Sim-to-Real experiments demonstrate that geometry from simulation can be faithfully combined with real-world appearance references. By decoupling geometry and appearance, our method improves robustness and generalization, requiring only diverse reference images rather than per-scene 3D scans or carefully aligned inputs, making it scalable for real-world deployment.
10.2 Scalability of 3D Data Acquisition
Globally consistent 3D conditions are essential for wrist-view camera motion. While acquiring dense, high-quality 3D data can be resource-intensive, our framework leverages automated, cost-efficient pipelines:
1. Real-robot data: RGB-D cameras capture metric depth alongside RGB images. For both RGB-D and RGB-only images, depth and normal maps are robustly generated using automated pipelines. The code will be released to ensure reproducibility.
2. Simulation data: Per-view 3D geometry conditions (depth and normals) are trivially available at no extra cost, making simulation a scalable source of 3D-aware inputs.
By combining real and simulated pipelines, our framework reduces the effective cost of 3D data acquisition, enabling high diversity and fine-grained controllability in synthetic video generation for robotic policy learning.
10.3 Rationale for Reference Selection
RoboTransfer is designed for scalable, fully automated training. Capturing a clean background for every demonstration would be prohibitively expensive. Instead, background imagery is extracted from existing videos, preserving fidelity without extra data collection. In inference, a clean background frame can be captured if desired, keeping the process simple and efficient. This design balances automation, scalability, and flexibility across training and deployment.